OpenStudy (m.d.saxon):

79*34*56a/34+25=115049 solve for a A.13 B.27 C.25 D.26

OpenStudy (johnweldon1993):

\[\large \frac{79\times 34\times 56a}{34 + 25} = 115049\] Solve the bottom first \[\large \frac{79\times 34\times 56a}{59} = 115049\] Multiply both sides of this equation by 59 \[\large 79\times 34\times 56a = 6787891\] Multiply everything \[\large 150416a = 6787891\] Divide both sides by 150416 \[\large a = 45.127\]

OpenStudy (m.d.saxon):

nope

OpenStudy (johnweldon1993):

How is the question written? As I've done it above? Or different?

OpenStudy (m.d.saxon):

the way I have it written

OpenStudy (m.d.saxon):

parenthesis exponent multiply divide add subtract

OpenStudy (johnweldon1993):

So written exactly how it is would be \[\large 79 \times 34 \times \frac{56a}{34} + 25 = 115049\] so if that's the case, subtract 25 from both sides \[\large 79 \cancel{\times 34} \times \frac{56a}{\cancel{34}} = 115024\] Multiply out 79 and 56 \[\large 4424a = 115024\] And divide both sides by 4424 \[\large a = 26\]

OpenStudy (m.d.saxon):

ok my mom saids that's right thanx I will metal you